# Sommerfeld Enhancement for Vector Dark Matter in the Fundamental Representation of $SU(2)_{L}$

### v1

@@ 3.4.1 Gauge Bosons-Dark Matter Interaction
We aim to investigate the annihilation of dark matter particles into standard model ones. Specifically, we consider the pairs ($v_{0}v_{0}$), for which such annihilation occurs. The most important contributions come from processes involving the exchange of $Z$ and $W^{\pm}$ bosons, having in their initial states ($v_{0}v_{0}$) , ($v_{0}a_{0}$), or ($v_{+}v_{-}$) pairs. Our approach involves deriving an effective action that accurately describes these pairs. For this purpose, we define the following vector:
@@
In the non-relativistic limit, the first term of equation (3.4.1) characterizes the Coulomb interaction between non-relativistic particles, whereas the remaining terms depict the Yukawa potentials related to the exchange of $Z$ and $W$ bosons. Upon employing vector (40), the effective action transforms into the following expression:
$S_{eff,1_{NR}}=-\int d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}V_{Gauge}(|\vec{x}-\vec{y}|)s(x,\vec{y})$ (42)
where $V_{Gauge}(|\vec{x}-\vec{y}|)$ is electroweak potential matrix (interaction gauge boson-vector dark matter), with $r=|\vec{x}-\vec{y}|$;
$V_{Gauge}(r)=-\dfrac{g^{2}}{8\pi r}\begin{pmatrix}4s_{W}^{2}e^{-M_{Z}r}&0&0&0\\ 0&4s_{W}^{2}e^{-M_{Z}r}&0&0\\ 0&0&16s_{W}^{2}e^{-M_{Z}r}&2\sqrt{2}e^{-M_{W}r}\\ 0&0&2\sqrt{2}e^{-M_{W}r}&8s_{W}^{2}e^{-M_{Z}r}-2(c_{W}-s_{W}t_{W}^{2})^{2}\end{pmatrix}$ (47)
In appendix A we show how to compute this matrix.
@@ 3.4.2 Higgs Boson-Dark Matter Interaction
We need to take into account the interaction between the Higgs Boson and the Dark Matter vector. Our primary focus lies in understanding the cubic interaction, as it is crucial for unraveling the nature of this interaction. For this purpose, we parametrize the Higgs doublet as folloes:
@@
where $\phi^{+}$ and $\phi^{0}$ are would-be Goldtone bosons, and $v$ is vaccum expectation value.
@@
$V_{h-V}(r)=\dfrac{v^{2}}{32\pi m^{2}r}\begin{pmatrix}8\lambda_{4}^{2}a&0&0&0\\ 0&8\lambda_{4}^{2}a&0&0\\ 0&0&\dfrac{1}{2}(\lambda_{2}+\lambda_{3})^{2}a&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}\\ 0&0&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}&\lambda_{2}^{2}a\end{pmatrix}$ (55)
In Appendix A, we illustrate the methodology for computing this matrix.
@@ 3.5.1 Gauge Boson-Dark Matter Vector
$\Gamma_{WW}^{J=0}=\dfrac{2g^{4}}{64\sqrt{3}\pi m^{2}}\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (61)
$\Gamma_{WW}^{J=2}=\dfrac{g^{4}}{64\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (66)
$\Gamma_{AA}^{J=0}=\dfrac{e^{4}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (71)
$\Gamma_{AA}^{J=2}=\dfrac{e^{4}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (76)
@@
$\Gamma_{ZZ}^{J=0}$ $=\dfrac{g^{4}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (81)
$\Gamma_{ZZ}^{J=2}$ $=\dfrac{g^{4}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (86)
@@
$\Gamma_{AZ}^{J=0}$ $=\dfrac{e^{2}g^{2}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (91)
$\Gamma_{AZ}^{J=2}$ $=\dfrac{e^{2}g^{2}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (96)
@@ 3.5.2 Higgs Boson-Dark Matter Vector
$\Gamma_{Higgs}=\dfrac{1}{32\pi m^{2}}\begin{pmatrix}\left(\dfrac{3}{2}+\sqrt{2}\right)^{2}\lambda_{4}^{2}&0&0&0\\ 0&\dfrac{9}{4}\lambda_{4}^{2}&0&0\\ 0&0&\dfrac{1}{4}(7\lambda_{2}^{2}+6\lambda_{2}\lambda_{3}+5\lambda_{3}^{2})&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}\\ 0&0&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}&\dfrac{3}{4}\lambda_{2}^{2}+(\lambda_{2}+\lambda_{3})^{2}\end{pmatrix}$ (101)
3.6 Results
@@
Our results are shown in figure 2 . There are several noteworthy characteristics of the Sommerfeld enhancement that merit discussion. Firstly, it is a velocity-dependent phenomenon. For instance, no structure is observed for a relative velocity of $v=0.3~{}c$ (dashed lines in figure 2), which is representative of the early Universe, when the relic density was formed. In contrast, for smaller velocities, like the one expected in the galaxy ($v=10^{-3}c$), significant oscillations and resonances appear. For masses that exhibit such resonances (roughly around $6$ TeV), the Sommerfeld factor experiences a significant boost of about 5 order of magnitude, for all the values of portal coupling constant $\lambda$.
Figure 2: Annihilation cross-section (fundamental representation spin-1 of $SU(2)_{L}$) for the process $v_{0}a_{0}\longrightarrow SMSM$ with the Sommerfeld Enhancement factor, considering velocities $v=10^{-3}$ (galactic Halo, shown as continuous lines, green, blue, black and gray lines ) and $v=0.3$ (Relic density, indicated by dashed lines), and different values of the Higgs portal coupling constant $\lambda$. The green line indicates that it is the spin-1 triplet representation of $SU(2)_{L}$, while the red line indicates that it is the spin-1/2 doublet representation of $SU(2)_{L}$.
@@ 4 Photons Flux from Annihilation neutral Dark Matter particles
$\dfrac{d\Phi}{dE}=\dfrac{1}{8\pi}J(\Omega)\dfrac{\braket{\sigma v}_{\gamma\gamma}}{m^{2}}\left(\dfrac{dN}{dE}\right)_{\gamma\gamma}$ (103)
@@
The $J$-factors for dwarf galaxies are approximately $J\approx 10^{19}$ $GeV^{2}/cm^{5}$, for the Andromeda galaxy, $J\approx 10^{20}$ $GeV^{2}/cm^{5}$, and for the Milky Way’s Galactic Center, $J\approx 10^{25}$ $GeV^{2}/cm^{5}$. If the solid angle changes by one degree, the $J$-factor undergoes an approximate change of $J\approx 10^{24}$ $GeV^{2}/cm^{5}$.
The distribution, denoted as $\left(\frac{dN}{dE}\right)_{\gamma\gamma}$, characterizes the number of photons with energy $E_{\gamma\gamma}=m$, where $m$ represents the mass of the WIMP. The distribution is defined as follows:
$\left(\dfrac{dN}{dE}\right)_{\gamma\gamma}=2\times\delta(E-E_{\gamma\gamma})$ (105)
while $\braket{\sigma v}_{\gamma\gamma}$ represents the annihilation cross-section rate with the mediator, which is dependent on the branching ratio, and is defined as $\braket{\sigma v}_{\gamma\gamma}=BR_{\gamma\gamma}\times\braket{\sigma v}_{XX^{\prime}}$. Here, $\braket{\sigma v}_{XX^{\prime}}$ corresponds to the total cross-section for Dark Matter annihilation, as shown in equation (30).
$BR_{\gamma\gamma}=\dfrac{\Gamma_{\gamma\gamma}}{\sum_{V}\Gamma_{VV^{\prime}}}$ (106)
where $V$ is the gauge and Higgs bosons. Branching ratio is the quotient between annihilation matrix of photons by add total annihilation matrix.
4.1 Results
Using the expressions above, we calculate the photon flux resulting from the Annihilation Cross Section (30), which serves as a method for the Indirect Detection of Dark Matter. This entails the integration of equation (103) using the distribution of the number of photons as defined in (105). Figure 3 depict the photon flux from VDM annihilation per year. We can see, as expected, that as the Higgs portal increases in size, the photon flux also increases.Our constructions are not constrained by Fermi-LAT, as our region of interest is above $100$ GeV which is not tested by Fermi-LAT; therefore, in the graphs, we compare with the projected CTA sensitivity [16]. We can see that for low vlues of the Higgs-DM coupling CTA can test only the region around the big resonance while for larger values CTA will be able to observe or rule out all our region of interest.
Figure 3: Photons Flux from Annihilation of neutral Dark Matter particles (Yellow, Dark Orange, Purple and Dark Pink lines correspond to fundamental representation of $SU(2)_{L}$ spin-1), black line is the projected CTA photon flux sensitivity, The green line corresponds to the photon flux of the spin-1 $SU(2)_{L}$ triplet representation, while the red line corresponds to the $SU(2)_{L}$ doublet representation for spin-1/2.
@@ 5 Summary
In this paper, we have studied the Sommerfeld enhancement in a model where the DM candidate is a neutral component of a new massive vector field in the fundamnetal representation of $SU(2)_{L}$. We found that big resonance in $\sigma v$ for DM masses in the rage of 5 TeV to 10 TeV, in the low velocities regime. We used this result to predict the photon flux produced by DM annihilation at the center of the galaxy. This prediction can be tested in the near future by CTA.

### v2

@@ 3.4.1 Gauge Bosons-Dark Matter Interaction
We aim to investigate the annihilation of dark matter particles into standard model ones. Specifically, we consider the pairs ($v_{0}v_{0}$), for which such annihilation occurs. The most important contributions come from processes involving the exchange of $\gamma$, $Z$, $W^{\pm}$ bosons, having in their initial states ($v_{0}v_{0}$) , ($v_{0}a_{0}$), or ($v_{+}v_{-}$) pairs. Our approach involves deriving an effective action that accurately describes these pairs. For this purpose, we define the following vector:
@@
In the non-relativistic limit, the first term of equation (3.4.1) characterizes the Coulomb potential between non-relativistic particles, whereas the remaining terms depict the Yukawa potentials related to the exchange of $Z$ and $W$ bosons. Upon employing vector (32), the effective action transforms into the following expression:
$S_{eff,1_{NR}}=-\int d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}V_{g-V}(|\vec{x}-\vec{y}|)s(x,\vec{y})$ (34)
where $V_{g-V}(|\vec{x}-\vec{y}|)$ is electroweak potential matrix (interaction gauge boson-vector dark matter), with $r=|\vec{x}-\vec{y}|$;
$V_{g-V}(r)=-\dfrac{g^{2}}{8\pi r}\begin{pmatrix}4a^{2}e^{-M_{Z}r}&0&0&0\\ 0&4a^{2}e^{-M_{Z}r}&0&0\\ 0&0&16a^{2}e^{-M_{Z}r}&b^{2}e^{-M_{W}r}\\ 0&0&b^{2}e^{-M_{W}r}&8a^{2}e^{-M_{Z}r}-2(c_{W}(1-(\kappa_{1}+\kappa{2})-s_{W}t_{W}^{2})^{2}\end{pmatrix}$ (35)
with $a=(c_{W}(\kappa_{1}+\kappa_{2})+s_{W})$, $b=(\frac{3}{8}-2\sqrt{2})(\kappa_{1}+\kappa_{2})+2\sqrt{2}$ In appendix A we show how to compute this matrix.
@@ 3.4.2 Higgs Boson-Dark Matter Interaction
We need to take into account the interaction between the Higgs Boson and the Dark Matter vector. Our primary focus lies in understanding the cubic interaction, as it is crucial for unraveling the nature of this interaction. For this purpose, we parameterize the Higgs doublet as follows:
@@
where $\phi^{+}$ and $\phi^{0}$ are would-be Goldtone bosons, and $v$ is vacuum expectation value.
@@
$V_{h-V}(r)=\dfrac{v^{2}}{32\pi M_{V}^{2}r}\begin{pmatrix}8\lambda_{4}^{2}a&0&0&0\\ 0&8\lambda_{4}^{2}a&0&0\\ 0&0&\dfrac{1}{2}(\lambda_{2}+\lambda_{3})^{2}a&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}\\ 0&0&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}&\lambda_{2}^{2}a\end{pmatrix}$ (37)
with $a=(2e^{-M_{h}r}+e^{-M_{Z}r})$. In Appendix A, we illustrate the methodology for computing this matrix.
@@ 3.5.1 Gauge Boson-Dark Matter Vector
$\Gamma_{WW}^{J=0}=\dfrac{2g^{4}}{64\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (39)
$\Gamma_{WW}^{J=2}=\dfrac{g^{4}}{64\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (40)
$\Gamma_{AA}^{J=0}=\dfrac{e^{4}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (41)
$\Gamma_{AA}^{J=2}=\dfrac{e^{4}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (42)
@@
$\Gamma_{ZZ}^{J=0}$ $=\dfrac{g^{4}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (43)
$\Gamma_{ZZ}^{J=2}$ $=\dfrac{g^{4}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$
@@
$\Gamma_{AZ}^{J=0}$ $=\dfrac{e^{2}g^{2}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (44)
$\Gamma_{AZ}^{J=2}$ $=\dfrac{e^{2}g^{2}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$
@@ 3.5.2 Higgs Boson-Dark Matter Vector
$\Gamma_{Higgs}=\dfrac{1}{32\pi M_{V}^{2}}\begin{pmatrix}\left(\dfrac{3}{2}+\sqrt{2}\right)^{2}\lambda_{4}^{2}&0&0&0\\ 0&\dfrac{9}{4}\lambda_{4}^{2}&0&0\\ 0&0&\dfrac{1}{4}(7\lambda_{2}^{2}+6\lambda_{2}\lambda_{3}+5\lambda_{3}^{2})&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}\\ 0&0&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}&\dfrac{3}{4}\lambda_{2}^{2}+(\lambda_{2}+\lambda_{3})^{2}\end{pmatrix}$ (45)
3.6 Numerical Results
@@
Our results are shown in figures 2 and 3 . The first one shows our predictions without taking into account the non-minimal interactions ($\kappa=0$) while in the second figure we used $\kappa=1$. In both cases, there are several noteworthy characteristics of the Sommerfeld enhancement that merit discussion. Firstly, as usual, the SE is a velocity-dependent phenomenon. For instance, no structure is observed for a relative velocity of $v=0.3$ (dashed lines in figures 2 and 3), which is representative of the early Universe, when the relic density was formed. In contrast, for smaller velocities, like the one expected in the galaxy ($v=10^{-3}$), significant oscillations and resonances appear. We have observed that the small oscillations dependent on interactions mediated by the Higgs boson. Indeed, we see that they are attenuated for large values of $\lambda$ parameter Interestingly, the presence of the non-minimal interactions introduce an additional resonance at a lower mass. However, the $\kappa$-terms don’t modify significantly the size of the oscillations nor the amplitude of the resonances. For masses that exhibit such resonances (roughly around $1.9$ and $6$ TeV), the Sommerfeld factor experiences a significant boost of about 5 order of magnitude, for all the values of portal coupling constant $\lambda$.
Figure 2: Annihilation cross-section (fundamental representation spin-1 of $SU(2)_{L}$) for the process $v_{0}a_{0}\longrightarrow SMSM$ with the Sommerfeld Enhancement factor, considering velocities $v=10^{-3}$ (galactic Halo, shown as continuous lines, green, blue, black and gray lines ) and $v=0.3$ (Relic density, indicated by dashed lines), and different values of the Higgs portal coupling constant $\lambda$, with $\kappa_{1}=\kappa_{2}=0$. The green line indicates that it is the spin-1 triplet representation of $SU(2)_{L}$, while the red line indicates that it is the spin-1/2 doublet representation of $SU(2)_{L}$.
Figure 3: Annihilation cross-section (fundamental representation spin-1 of $SU(2)_{L}$) for the process $v_{0}a_{0}\longrightarrow SMSM$ with the Sommerfeld Enhancement factor, considering velocities $v=10^{-3}$ (galactic Halo, shown as continuous lines, green, blue, black and gray lines ) and $v=0.3$ (Relic density, indicated by dashed lines), and different values of the Higgs portal coupling constant $\lambda$, with $\kappa_{1}=\kappa_{2}=1$. The green line indicates that it is the spin-1 triplet representation of $SU(2)_{L}$, while the red line indicates that it is the spin-1/2 doublet representation of $SU(2)_{L}$.
@@ 4 Photons Flux from Annihilation neutral Dark Matter particles
$\dfrac{d\Phi}{dE_{\gamma}}=\sum_{i}\dfrac{1}{8\pi}J(\Omega)\dfrac{\braket{\sigma v}_{i}}{M_{V}^{2}}\left(\dfrac{dN_{i}}{dE_{\gamma}}\right)$ (47)
@@
where $\rho[r(\ell,\psi)]$ is the DM density and depends of the assumed astrophysical profile. The $J$-factor for dwarf galaxies is approximately $J\approx 10^{19}$ GeV2/cm5 while for the Andromeda galaxy, $J\approx 10^{20}$ GeV2/cm5, and for the Milky Way’s Galactic Center, $J\approx 10^{25}$ GeV2/cm5. If the solid angle changes by one degree, the $J$-factor undergoes an approximate change of $J\approx 10^{24}$ GeV2/cm5 [20].
The distribution, denoted as $\left(\frac{dN_{i}}{dE_{\gamma}}\right)$, characterizes the number of photons with energy $E_{\gamma}$, while $\braket{\sigma v}_{i}$ represents the averaged annihilation cross-section from the i-th channel of annihilation. This quantity dependents on the branching ratio, and is defined as $\braket{\sigma v}_{\gamma\gamma}=BR_{\gamma\gamma}\times\braket{\sigma v}_{XX^{\prime}}$. Here, $\braket{\sigma v}_{XX^{\prime}}$ corresponds to the total cross-section for Dark Matter annihilation, as shown in equation (26).
Using the expressions above, we calculate the photon flux resulting from the annihilation cross section (26), which serves as a method for the indirect detection of DM. This entails the integration of equation (47). Figures 4 and 5 depict the photon flux from VDM annihilation per second for $\kappa=0$ and $\kappa=1$ respectively. In all the cases we have used $\lambda=0.5$. We present the signals of photon flux resulting from the annihilation of VDM particles from the center of the galaxy for different final states ($\gamma\gamma$, $W^{+}W^{-}$, $ZZ$, and $hh$), with a solid angle of $1^{\circ}$, compared with the canonical cross-section with a mass of 100 GeV. Additionally, we show the total flux predicted at different values of the DM mass. We selected the masses to coincide with peaks and local minimums in the SE factor. We compare our results with Fermi-LAT sesitivity and the projected CTA sensitivity [21].
Figure 4: Photons Flux from Annihilation of neutral Dark Matter particles. Where it is done for the Navarro-Frenk-White (NFW) profile, with a Line of sight (los) angle in longitude (llos) of $1^{\circ}$, and latitude (blos) of $0^{\circ}$. This photon flux is measured at the center of the galaxy, with $\kappa=0$. We show differential sensitivity of CTA [21] and Fermi-Lat [22] in the galactic center.
Figure 5: Photons Flux from Annihilation of neutral Dark Matter particles. Where it is done for the Navarro-Frenk-White (NFW) profile, with a Line of sight (los) angle in longitude (llos) of $1^{\circ}$, and latitude (blos) of $0^{\circ}$. This photon flux is measured at the center of the galaxy, with $\kappa=1$. We show differential sensitivity of CTA [21] and Fermi-Lat [22] in the galactic center.
@@ 5 Summary
In this paper, we have studied the Sommerfeld enhancement in a model where the DM candidate is a neutral component of a new massive vector field in the fundamental representation of $SU(2)_{L}$. We found that the annihilation cross-section predicted by our model is generally bigger than the one predicted in the cases of a spin $1/2$ doublet and a spin $1$ triplet. Additionaly, the presence of non-minimal interactions of the doublet with the standard gauge model produces a richer resonance structure. In fact, in the absence of the non-minimal interaction, only one big resonance in $\sigma v$ is observed for $M_{V}\approx 6$ TeV. However, when the non-minimal interactions are present, a second peak appears at $M_{V}\approx 2$ TeV, in the low velocities regime.
As a product of our work, we used our results on $\sigma v$ to predict the photon flux produced by DM annihilation at the center of the galaxy.